# UniLGL: Learning Uniform Place Recognition for FOV-limited/Panoramic LiDAR Global Localization

### v1

@@ V-B1 Ablation Study
| Sequence | UniLGL | UniLGL w/o intensity | UniLGL w/o spatial | UniLGL w/o loc. feat. | BEVPlace++ | Logg3D-Net | RING++ | |
| Fov-limited LiDAR | Mid_NTU_02 | 83.92/1.12/0.59 | 80.47/1.21/0.60 | 80.25/1.11/0.69 | 35.44/1.49/0.71 | 55.19/1.94/0.57 | 62.13/1.08/0.38 | 31.02/2.07/0.85 |
| Mid_NTU_10 | 91.06/0.66/0.35 | 90.44/0.70/0.41 | 87.77/0.67/0.45 | 80.07/0.98/0.44 | 88.70/1.10/0.31 | 93.75/0.78/0.28 | 57.08/1.45/0.78 | |
| Mid_NTU_13 | 80.76/1.31/0.59 | 77.57/1.37/0.62 | 78.32/1.28/0.68 | 42.91/1.29/0.52 | 55.31/1.77/0.51 | 58.15/1.33/0.41 | 32.74/1.65/0.81 | |
| Garden_01 | 91.44/0.63/0.28 | 90.64/0.65/0.33 | 89.17/0.63/0.32 | 80.42/0.71/0.32 | 91.59/0.68/0.18 | 91.77/0.63/0.31 | 83.33/1.12/0.67 | |
| Garden_02 | 91.58/0.70/0.32 | 90.59/0.70/0.35 | 91.00/0.73/0.36 | 73.62/0.86/0.38 | 89.05/0.78/0.20 | 91.47/0.76/0.34 | 83.42/1.22/0.68 | |
| Garden_03 | 90.55/0.71/0.31 | 86.70/0.73/0.34 | 87.15/0.72/0.37 | 69.35/0.91/0.41 | 85.35/0.79/0.20 | 87.58/0.71/0.33 | 67.12/1.24/0.73 | |
| Garden_04 | 95.34/0.71/0.32 | 95.16/0.77/0.35 | 88.24/0.72/0.34 | 76.87/0.83/0.35 | 90.00/0.77/0.20 | 95.79/0.71/0.33 | 74.23/1.23/0.74 | |
| Average | 89.24/0.83/0.39 | 87.37/0.88/0.43 | 85.99/0.84/0.46 | 65.53/1.01/0.45 | 79.31/1.12/0.31 | 82.95/0.86/0.34 | 61.28/1.43/0.75 | |
| Panoramic LiDAR | OS_NTU_02 | 99.82/0.69/0.27 | 99.74/0.66/0.28 | 98.45/0.55/0.45 | 63.93/1.10/0.41 | 96.15/1.97/0.42 | 86.56/0.79/0.27 | 75.51/1.44/0.50 |
| OS_NTU_10 | 99.66/0.37/0.18 | 99.65/0.38/0.19 | 98.88/0.36/0.26 | 96.37/0.51/0.22 | 99.38/1.10/0.17 | 99.07/0.43/0.16 | 75.57/1.44/0.50 | |
| OS_NTU_13 | 100.0/1.06/0.44 | 99.86/1.08/0.48 | 99.87/1.05/0.54 | 60.90/1.53/0.60 | 96.33/1.86/0.52 | 78.15/1.50/0.50 | 58.65/1.91/0.60 | |
| Snail_81R_02 | 79.81/0.83/1.04 | 78.60/0.82/1.07 | 63.30/1.03/0.98 | 11.79/1.82/1.11 | 73.06/1.27/1.06 | 31.51/2.07/0.90 | 29.88/1.40/0.94 | |
| Snail_81R_03 | 81.82/0.75/0.50 | 82.70/0.79/0.57 | 82.19/0.75/0.63 | 14.51/1.89/0.90 | 80.39/1.31/0.65 | 40.57/2.03/0.64 | 31.46/1.52/0.61 | |
| Garden_LT_01 | 81.13/1.21/0.52 | 64.24/1.36/0.57 | 56.39/1.37/0.71 | 1.69/2.20/0.75 | 13.17/1.99/0.77 | 19.75/1.80/0.52 | 77.23/1.65/0.48 | |
| Garden_LT_02 | 75.16/1.24/0.67 | 61.93/1.28/0.55 | 51.89/1.41/0.74 | 1.08/1.88/0.71 | 17.46/1.29/0.77 | 20.20/1.91/0.56 | 61.35/1.62/0.47 | |
| Garden_LT_03 | 76.28/1.30/0.53 | 71.62/1.46/0.56 | 55.40/1.63/0.68 | 4.38/2.31/0.87 | 18.07/1.23/0.75 | 26.40/1.81/0.56 | 53.21/1.58/0.44 | |
| Garden_LT_04 | 79.39/1.30/0.56 | 77.42/1.39/0.57 | 57.81/1.67/0.67 | 5.32/1.68/0.89 | 18.25/1.30/0.79 | 21.32/1.57/0.55 | 61.61/1.60/0.48 | |
@@
For the results shown in Table. III, the proposed UniLGL achieves an average success rate of and when using FoV-limited LiDAR and panoramic LiDAR, respectively. To validate the effectiveness of introducing spatial and material uniformity, we compare UniLGL with two ablated variants, UniLGL w/o Intensity and UniLGL w/o Spatial. The proposed UniLGL achieves a – and – improvement in global localization success rate compared to the respective variants that exclude material and spatial uniformity. For challenging sequences with long time intervals, such as Garden_LT_01 to Garden_LT_04, UniLGL achieves more robust global localization performance, with average improvements of and in success rate over UniLGL w/o Intensity and UniLGL w/o Spatial, respectively. For large-scale urban driving sequences such as Snail_81R_02 and Snail_81R_03, UniLGL w/o Intensity and UniLGL w/o Spatial achieve comparable global localization success rates to UniLGL on Snail_81R_03, but are outperformed by UniLGL on Snail_81R_02. This discrepancy can be attributed to the difference in dynamic object density, where Snail_81R_02 was recorded during peak traffic hours, while Snail_81R_03 was collected at night during off-peak hours. These results indicate that UniLGL offers a more robust global localization to dynamic objects, such as moving vehicles and pedestrians, compared to the ablated variants. To illustrate the effectiveness of introducing viewpoint invariance at the local feature level, an ablation study is also conducted by comparing models with and without viewpoint-invariant local feature supervision. From the comparison result, UniLGL consistently outperforms UniLGL w/o Loc. Feat. across all 16 sequences, achieving an average improvement of – in global localization success rate on both FoV-limited and panoramic LiDAR. For metric localization, UniLGL fuses pixel-level viewpoint-invariant local features from spatial and intensity BEV images and maps them to the point cloud, further delivering more accurate 6-DoF metric localization compared with its ablated variants.
@@ VI-B Application 2: Multi-MAV Collaborative Exploration with Bandwidth Limitation
To further validate its practicality, UniLGL is deployed on a multi-MAV system constructed by 4 identical MAVs to enable collaborative localization during an exploration task in a field scenario. As shown in Fig. 8(b), each MAV platform is equipped with a Livox Mid 360 LiDAR, a Pixhawk4 flight controller, a low-power onboard computer (Nvidia Orin NX), and a VEWOE VMA10A mesh networking unit. During collaborative exploration, we integrate the observations from UniLGL and CTE-MLO into a factor graph to enable decentralized collaborative state estimation for multiple MAVs. As illustrated in Fig. 10, the relative pose estimations provided by UniLGL are incorporated as inter-MAV measurements, while the odometry information from CTE-MLO is utilized as intra-MAV measurements within the factor graph. As illustrated in Fig. 11(a), UniLGL provides sufficient relative localization constraints among multiple MAVs, enabling simultaneous infrastructure-free collaborative localization and mapping. Notably, the use of UniLGL for inter-MAV constraints significantly reduces the communication bandwidth required in multi-agent systems. This efficiency is reflected in two main aspects: 1) UniLGL enables collaborative localization through an event-triggered mechanism, where a UAV transmits relative localization information to others only after a successful global descriptor matching, rather than continuously broadcasting point clouds as required by methods such as [77, 78, 79]. 2) For relative pose estimation, UniLGL requires transmitting only a small number of point cloud keypoints associated with matched local features, thereby eliminating the need to transmit raw point clouds for pose refinement. As shown in Fig. 11(c), the UniLGL-based collaborative localization approach achieves substantially lower communication bandwidth consumption compared to methods that lack event-triggered mechanisms (noted as Point Cloud in Fig. 11(c)) or require additional point cloud registration (noted as Point Cloud. Event in Fig. 11(c)). Owing to the superior efficiency of UniLGL in terms of computational and communication resource consumption, we integrate UniLGL-based collaborative localization with a decentralized exploration algorithm DPPM[80] and the MAV trajectory tracking control algorithm FxTDO-MPC[81], which enables a real-world exploration fully onboard. As shown in Fig. 11(c), the multi-MAV system completed high-precision scanning in a dense forest environment. The area of the exploration regions reaches .
@@ VII Conclusion
In this article, a Uniform LiDAR-based Global Localization system, UniLGL, is developed to achieve cascaded place recognition and global pose estimation with the consideration of spatial and material uniformity as while as sensor-type uniformity. To equip UniLGL with spatial and material uniformity, we represent the 4D point cloud information in a lossless manner using an image pair consisting of a spatial BEV image and an intensity BEV image, and design an end-to-end BEV fusion network for place recognition. For sensor-type uniformity, a viewpoint invariance hypothesis is introduced to replace the conventional translation equivariance assumption commonly used in existing LPR networks[24, 15, 14, 51, 25], which hypothesis guides UniLGL to learn global descriptors and local features with consistency across geographically distant but co-visible areas (as demonstrated in Fig. 3). Moreover, a vision foundation model, DINO[63], is elegantly integrated into the proposed BEV fusion network to enhance the generalization capability of global descriptors and local features, without requiring large amounts of LiDAR training data. Thanks to the consistency of local features across co-visible areas, a global pose estimator is derived using graduated non-convexity optimization[69] to estimate the 6-DoF global pose based on point-level correspondences established through local feature matching between BEV images. With the successful integration of a series of theory and practical implementations (e.g. uniform LPR, task agnostic vision foundation model, and robust global estimator on SE(3)) into the UniLGL, the proposed method is demonstratively competitive compared to SOTA LPR/LGL methods. The extensive experimental results have demonstrated that the proposed UniLGL remains robust under extremely challenging conditions, such as long-term global localization and large viewpoint variations (e.g., reverse direction, large translation), across heterogeneous LiDAR configurations. Furthermore, the proposed UniLGL has been deployed to support autonomy on diverse platforms, from full-size trucks to power-limited MAV, which enables high-precision truck localization and mapping in a port environment and multi-MAV collaborative exploration in a forest environment. These real-world deployments affirm the extendability and applicability of UniLGL in industrial and field scenarios.

### v2

@@ V-B1 Ablation Study
Sequence | UniLGL |
BEVPlace++ | Logg3D-Net | RING++ | |||||||
FoV-limited LiDAR |
Mid_NTU_02 | 83.92/1.12/0.59 | 80.47/1.21/0.60 | 80.25/1.11/0.69 | 35.44/1.49/0.71 | 55.19/1.94/0.57 | 62.13/1.08/0.38 | 31.02/2.07/0.85 | ||||||
Mid_NTU_10 | 91.06/0.66/0.35 | 90.44/0.70/0.41 | 87.77/0.67/0.45 | 80.07/0.98/0.44 | 88.70/1.10/0.31 | 93.75/0.78/0.28 | 57.08/1.45/0.78 | |||||||
Mid_NTU_13 | 80.76/1.31/0.59 | 77.57/1.37/0.62 | 78.32/1.28/0.68 | 42.91/1.29/0.52 | 55.31/1.77/0.51 | 58.15/1.33/0.41 | 32.74/1.65/0.81 | |||||||
Garden_01 | 91.44/0.63/0.28 | 90.64/0.65/0.33 | 89.17/0.63/0.32 | 80.42/0.71/0.32 | 91.59/0.68/0.18 | 91.77/0.63/0.31 | 83.33/1.12/0.67 | |||||||
Garden_02 | 91.58/0.70/0.32 | 90.59/0.70/0.35 | 91.00/0.73/0.36 | 73.62/0.86/0.38 | 89.05/0.78/0.20 | 91.47/0.76/0.34 | 83.42/1.22/0.68 | |||||||
Garden_03 | 90.55/0.71/0.31 | 86.70/0.73/0.34 | 87.15/0.72/0.37 | 69.35/0.91/0.41 | 85.35/0.79/0.20 | 87.58/0.71/0.33 | 67.12/1.24/0.73 | |||||||
Garden_04 | 95.34/0.71/0.32 | 95.16/0.77/0.35 | 88.24/0.72/0.34 | 76.87/0.83/0.35 | 90.00/0.77/0.20 | 95.79/0.71/0.33 | 74.23/1.23/0.74 | |||||||
Average | 89.24/0.83/0.39 | 87.37/0.88/0.43 | 85.99/0.84/0.46 | 65.53/1.01/0.45 | 79.31/1.12/0.31 | 82.95/0.86/0.34 | 61.28/1.43/0.75 | |||||||
Panoramic LiDAR |
OS_NTU_02 | 99.82/0.69/0.27 | 99.74/0.66/0.28 | 98.45/0.55/0.45 | 63.93/1.10/0.41 | 96.15/1.97/0.42 | 86.56/0.79/0.27 | 75.51/1.44/0.50 | ||||||
OS_NTU_10 | 99.66/0.37/0.18 | 99.65/0.38/0.19 | 98.88/0.36/0.26 | 96.37/0.51/0.22 | 99.38/1.10/0.17 | 99.07/0.43/0.16 | 75.57/1.44/0.50 | |||||||
OS_NTU_13 | 100.0/1.06/0.44 | 99.86/1.08/0.48 | 99.87/1.05/0.54 | 60.90/1.53/0.60 | 96.33/1.86/0.52 | 78.15/1.50/0.50 | 58.65/1.91/0.60 | |||||||
Snail_81R_02 | 79.81/0.83/1.04 | 78.60/0.82/1.07 | 63.30/1.03/0.98 | 11.79/1.82/1.11 | 73.06/1.27/1.06 | 31.51/2.07/0.90 | 29.88/1.40/0.94 | |||||||
Snail_81R_03 | 81.82/0.75/0.50 | 82.70/0.79/0.57 | 82.19/0.75/0.63 | 14.51/1.89/0.90 | 80.39/1.31/0.65 | 40.57/2.03/0.64 | 31.46/1.52/0.61 | |||||||
Garden_LT_01 | 81.13/1.21/0.52 | 64.24/1.36/0.57 | 56.39/1.37/0.71 | 1.69/2.20/0.75 | 13.17/1.99/0.77 | 19.75/1.80/0.52 | 77.23/1.65/0.48 | |||||||
Garden_LT_02 | 75.16/1.24/0.67 | 61.93/1.28/0.55 | 51.89/1.41/0.74 | 1.08/1.88/0.71 | 17.46/1.29/0.77 | 20.20/1.91/0.56 | 61.35/1.62/0.47 | |||||||
Garden_LT_03 | 76.28/1.30/0.53 | 71.62/1.46/0.56 | 55.40/1.63/0.68 | 4.38/2.31/0.87 | 18.07/1.23/0.75 | 26.40/1.81/0.56 | 53.21/1.58/0.44 | |||||||
Garden_LT_04 | 79.39/1.30/0.56 | 77.42/1.39/0.57 | 57.81/1.67/0.67 | 5.32/1.68/0.89 | 18.25/1.30/0.79 | 21.32/1.57/0.55 | 61.61/1.60/0.48 | |||||||
@@
For the results shown in Table. III, the proposed UniLGL achieves an average success rate of and when using FoV-limited LiDAR and panoramic LiDAR, respectively. To validate the effectiveness of introducing spatial and material uniformity, we compare UniLGL with two ablated variants, UniLGL w/o Intensity and UniLGL w/o Spatial. The proposed UniLGL achieves a – and – improvement in global localization success rate compared to the respective variants that exclude material and spatial uniformity on sequences collected by FoV-limited LiDAR and panoramic LiDAR, respectively. For challenging sequences with long time intervals, such as Garden_LT_01 to Garden_LT_04, UniLGL achieves more robust global localization performance, with average improvements of and in success rate over UniLGL w/o Intensity and UniLGL w/o Spatial, respectively. For large-scale urban driving sequences such as Snail_81R_02 and Snail_81R_03, UniLGL w/o Intensity and UniLGL w/o Spatial achieve comparable global localization success rates to UniLGL on Snail_81R_03, but are outperformed by UniLGL on Snail_81R_02. This discrepancy can be attributed to the difference in dynamic object density, where Snail_81R_02 was recorded during peak traffic hours, while Snail_81R_03 was collected at night during off-peak hours. These results indicate that UniLGL offers a more robust global localization to dynamic objects, such as moving vehicles and pedestrians, compared to the ablated variants. To illustrate the effectiveness of introducing viewpoint invariance at the local feature level, an ablation study is also conducted by comparing models with and without viewpoint-invariant local feature supervision. From the comparison result, UniLGL consistently outperforms UniLGL w/o Loc. Feat. across all 16 sequences, achieving an average improvement of – in global localization success rate on both FoV-limited and panoramic LiDAR. For metric localization, UniLGL fuses pixel-level viewpoint-invariant local features from spatial and intensity BEV images and maps them to the point cloud, further delivering more accurate 6-DoF metric localization compared with its ablated variants.
@@ VI-B Application 2: Multi-MAV Collaborative Exploration with Bandwidth Limitation
To further validate its practicality on lightweight power-limited platforms, UniLGL is deployed on a multi-MAV system constructed by 4 identical MAVs to enable collaborative localization during an exploration task in a field scenario. As shown in Fig. 9(b), each MAV platform is equipped with a Livox Mid 360 LiDAR, a Pixhawk4 flight controller, a low-power onboard computer (Nvidia Orin NX), and a VEWOE VMA10A mesh networking unit. During collaborative exploration, we integrate the observations from UniLGL and CTE-MLO into a factor graph to enable decentralized collaborative state estimation for multiple MAVs. As illustrated in Fig. 11, the relative pose estimations provided by UniLGL are incorporated as inter-MAV measurements, while the odometry information from CTE-MLO is utilized as intra-MAV measurements within the factor graph. As illustrated in Fig. 12(a), UniLGL provides sufficient relative localization constraints among multiple MAVs, enabling simultaneous infrastructure-free collaborative localization and mapping. Notably, the use of UniLGL for inter-MAV constraints significantly reduces the communication bandwidth required in multi-agent systems. This efficiency is reflected in two main aspects: 1) UniLGL enables collaborative localization through an event-triggered mechanism, where a UAV transmits relative localization information to others only after a successful global descriptor matching, rather than continuously broadcasting point clouds as required by methods such as [81, 82, 83]. 2) For relative pose estimation, UniLGL requires transmitting only a small number of point cloud keypoints associated with matched local features, thereby eliminating the need to transmit raw point clouds for pose refinement. As shown in Fig. 12(b), the UniLGL-based collaborative localization approach achieves substantially lower communication bandwidth consumption compared to methods that lack event-triggered mechanisms (noted as Point Cloud in Fig. 12(b)) or require additional point cloud registration (noted as Point Cloud. Event in Fig. 12(b)). Owing to the superior efficiency of UniLGL in terms of computational and communication resource consumption, we integrate UniLGL-based collaborative localization with a decentralized exploration algorithm DPPM[84] and the MAV trajectory tracking control algorithm FxTDO-MPC[85], which enables a real-world exploration fully onboard. As shown in Fig. 12(c), the multi-MAV system completed high-precision scanning in a dense forest environment. The area of the exploration regions reaches .
@@ VII Conclusion
In this article, a Uniform LiDAR-based Global Localization system, UniLGL, is developed to achieve cascaded place recognition and global pose estimation with the consideration of spatial and material uniformity as while as sensor-type uniformity. To equip UniLGL with spatial and material uniformity, we represent the 4D point cloud information in a lossless manner using an image pair consisting of a spatial BEV image and an intensity BEV image, and design an end-to-end BEV fusion network for place recognition. For sensor-type uniformity, a viewpoint invariance hypothesis is introduced to replace the conventional translation equivariance assumption commonly used in existing LPR networks[24, 15, 14, 54, 25], which hypothesis guides UniLGL to learn global descriptors and local features with consistency across geographically distant but co-visible areas (as demonstrated in Fig. 4). Moreover, a vision foundation model, DINO[68], is elegantly integrated into the proposed BEV fusion network to enhance the generalization capability of global descriptors and local features, without requiring large amounts of LiDAR training data. Thanks to the consistency of local features across co-visible areas, a global pose estimator is derived using graduated non-convexity optimization[74] to estimate the 6-DoF global pose based on point-level correspondences established through local feature matching between BEV images. With the successful integration of a series of theory and practical implementations (e.g. uniform LPR, task agnostic vision foundation model, and robust global estimator on SE(3)) into the UniLGL, the proposed method is demonstratively competitive compared to SOTA LPR/LGL methods. The extensive experimental results have demonstrated that the proposed UniLGL remains robust under extremely challenging conditions, such as long-term global localization and large viewpoint variations, across heterogeneous LiDAR configurations. Furthermore, the proposed UniLGL has been deployed to support autonomy on diverse platforms, from full-size trucks to power-limited MAV, which enables high-precision truck localization and mapping in a port environment and multi-MAV collaborative exploration in a forest environment. These real-world deployments affirm the extendability and applicability of UniLGL in industrial and field scenarios.
Appendix A Effectiveness of introduce Foundation Model into LGL & Zero-Shot Generalization across FoV-Limited and Panoramic LiDAR
Foundation models in robotics aim to provide systems with broad generalization capabilities by leveraging large-scale pretraining on diverse data sources. To facilitate understanding of the role of introducing foundation models into LGL, we evaluate the LGL performance of three variants: UniLGL, UniLGL initialized with a foundation model but without fine-tuning (UniLGL w/o FT), and UniLGL trained from scratch without foundation model initialization (UniLGL w/o FM). It is worth noting that, to demonstrate the zero-shot generalization ability, no cross-modal training is performed. We perform benchmark experiments on two datasets, MCD[76] and Garden[78], both of which contain paired FoV-limited and panoramic LiDAR measurements. During the experiments, the FoV-limited LiDAR scans are used as queries, while the panoramic LiDAR scans serve as the database, allowing us to assess the cross-modal generalization ability of LGL methods. The associated data sequences are referred to as NTU_CM_XX and Garden_CM_XX, respectively.
Sequence | UniLGL |
NTU_CM_02 | 97.75/88.63 | 37.95/16.38 | 0.60/4.54 | ||
NTU_CM_10 | 92.95/83.45 | 32.61/14.76 | 3.64/5.18 | ||
NTU_CM_13 | 96.46/88.49 | 41.02/17.50 | 8.59/7.19 | ||
Garden_CM_01 | 93.19/80.62 | 26.02/15.08 | 3.33/14.84 | ||
Garden_CM_02 | 91.86/84.73 | 30.54/19.04 | 17.54/18.97 | ||
Garden_CM_03 | 94.37/81.45 | 31.32/16.65 | 19.71/15.52 | ||
Garden_CM_04 | 87.37/86.03 | 37.06/19.82 | 24.86/15.59 | ||
Average | 93.42/85.20 | 33.78/17.03 | 11.18/11.69 |
1
The best result is highlighted in Bold.
Sequence | UniLGL |
NTU_CM_02 | 86.75/0.76/1.47 | 0.07/0.78/2.34 | 0.01/1.92/3.45 | ||
NTU_CM_10 | 78.19/0.65/1.34 | 0.13/1.46/3.44 | 0.01/1.54/3.92 | ||
NTU_CM_13 | 86.13/0.69/1.49 | 0.35/1.34/3.57 | - | ||
Garden_CM_01 | 75.77/0.86/2.44 | 0.25/1.44/3.54 | - | ||
Garden_CM_02 | 76.10/0.75/2.44 | 0.34/1.22/3.27 | 0.01/1.77/4.01 | ||
Garden_CM_03 | 76.07/0.77/2.53 | 0.41/1.19/2.58 | - | ||
Garden_CM_04 | 76.94/0.67/2.37 | 0.36/1.08/2.44 | 0.01/1.27/2.89 | ||
Average | 79.42/0.74/2.01 | 0.27/1.21/3.03 | 0.01/1.63/3.57 |
1
The best result is highlighted in Bold.
2
- denotes no successful global localization achieved.
As shown in the t-SNE[61] visualization in Fig. 13, introducing a foundation model imparts a certain level of generalization ability to the LPR network, enabling UniLGL w/o FT to achieve better clustering performance than UniLGL w/o FM, without fine-tuning. By initializing the network weights with DINO[68] and fine-tuning with only a small amount of homogeneous LiDAR data, UniLGL learns highly discriminative global descriptors in the heterogeneous LPR task. In addition to the above qualitative analysis, we present quantitative results of place recognition and global localization performance in Table V and Table VI, respectively. The results show that, thanks to the introduction of the foundation model, UniLGL achieves outstanding zero-shot cross-modal generalization ability. When compared to UniLGL w/o FT and UniLGL w/o FM, UniLGL achieves a - improvement in recall and a - improvement in average precision. For global localization, UniLGL achieves over increase in successful localization rate compared to UniLGL w/o FT and UniLGL w/o FM, with average global localization accuracy reaching and .